OpenStudy (anonymous):

help please !!! simplify completley:

OpenStudy (anonymous):

\[\frac{ t ^{4}-1 }{ t ^{3}+t ^{2}+t+1 }\] bottom equation is t^3 + t^2 + t + 1

OpenStudy (dinnertable):

Do you see anything that can be factored in the numerator?

OpenStudy (anonymous):

can it be \[{( t ^{2}-1)(t ^{2}+1) }\]

OpenStudy (dinnertable):

Yes it can

OpenStudy (dinnertable):

Do you see another application of the same thing within the numerator?

OpenStudy (anonymous):

no, i do not. can you please help me

OpenStudy (dinnertable):

Alright, so the first time you used difference of squares to factor \[t^4-1 = (t^2-1)(t^2+1)\]If you use difference of squares again, on the first factor(t^2-1), you will get a factored version of the numerator.\[t^4-1 = (t^2-1)(t^2+1) = (t-1)(t+1)(t^2+1)\]From here, you just need to factor the denominator so that you can cancel factors and simplify.

OpenStudy (dinnertable):

Do you see any way of factoring the denominator?

OpenStudy (anonymous):

no, i do not know how to factor four "numbers".

OpenStudy (dinnertable):

lol, have you learned about the factor theorem?

OpenStudy (dinnertable):

or synthetic divison for that matter?

OpenStudy (anonymous):

not that i recall, no

OpenStudy (dinnertable):

hmm, well the factor that I noticed was when I factored t^2 from the first two terms.\[t^3 + t^2 + t + 1 = t^2(t+1) + t+ 1\]What might be more noticeable now is that factoring by grouping is an option now.\[ t^2(t+1) + (t+1) = (t^2+1)(t+1)\]

OpenStudy (anonymous):

so the answer would be t-1 because the t^2+1 cancel out from both top and bottom and the t+1 cancel out as well, so whats left is t-1.

OpenStudy (dinnertable):

exactly

OpenStudy (anonymous):

okay, thank you so much for your help and time. i really appreciate it!(:

OpenStudy (dinnertable):

np